a potentiometer wire have a resistivity of 10^9 ohm-cm and area of cross section is 10^(-2) cm^2. if a current of 0.01mA passes through the wire. potential gradient is


1)10^9 v/m


2) 10^10 v/m


3) 10^11 v/m


4) 10^8 v/m

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

R=pl/A,V=IR=plI/A,potential gradient=V/l=pI/A=10^8 V/m